OpenStudy (anonymous):

Andrea's sleeping bag is designed for camping in temperatures of -9°C or warmer. For what Fahrenheit temperatures is her sleeping bag designed? Round to the nearest tenth of a degree. A. -22.8° F or warmer B. 41.4° F or warmer C. -37.0° F or warmer D. 15.8° F or warmer help needed plz I put c and it was worng. will give medals

OpenStudy (texaschic101):

formula for converting Celsius to Fahrenheit is : C x 9/5 + 32 = F so we have C = -9 now we sub -9 x 9/5 + 32 = F -81/5 + 32 = F -81/5 + 160/5 = F 79/5 = F 15.8 = F (answer is D) Remember this : C to F : multiply by 9, divide by 5, then add 32 F to C : deduct 32, multiply by 5, then divide by 9